Question 6
6. A three-phase, 50 Hz, 667 MVA, 22 kV, 4-pole steam driven turbine
generator has an
1
Expert's answer
2021-08-16T02:19:32-0400

The synchronous electrical radial frequency 

"\omega_{syn}= 2 \pi f \\
\omega_{syn}= 2 \pi (50)\\
\omega_{syn}= 314.2 rad/s"


The kinetic energy of rotating mass at the synchronous speed

"KE= HS_{rated}\\
KE= 4*667 *10^6\\
KE= 2668*10^6\\
KE=2.668*10^9 J"